OpenStudy (anonymous):

How do I solve this quadratic equation by factoring: 2x^2 - 3x = 0

OpenStudy (anonymous):

You need to take out the common factor in both terms. Can you see what it is?

OpenStudy (anonymous):

Isn't it product 2 and sum: -3?

OpenStudy (anonymous):

okay we have 2x^2 - 3x = 0 There is a x in the 3x term and a x in the 2x^2 term. THis is all we can take out.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

\[2x^2-3x=0\]\[x(2x-3)=0\] So x=0 Or 2x-3=0 2x=3 x=3/2

OpenStudy (anonymous):

ohhh.. so it's not like (2x^2 -2) + (-3x +1)...? thank you!!

OpenStudy (anonymous):

No worries,